Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day03


Crew 102 Log Book for March 28, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: Excellent

Time Departed/Returned from EVA: 2 EVAs 10:30 - 11:15, 11:45 - 13:00

Brief Narrative of Field Mission Results:

Inspection of the science project conducted on the previous days Inspection of the habitat’s structural corrosion

Collection of rock samples to check for cyanobacteria existence


EVA Data/Interpretations:

Samples and data analysis ongoing

Engineering/Hab Maintenance:

It seems the Water Heater leaking problem is getting worse. Being monitored.

We are currently try to save propane (low level, 14%).

Gray water system not functional at the moment. 

Plans for Tomorrow:

Careful maintenance of ETVs (oil, tires, etc.). The 3 ETVs in best shape will support a longer EVA aimed at the collection of more rock samples for the cyanobacteria study.

Restoring of Radiotelescope antenna

Continue on current science projects (psychological, physiological, Hub Monitoring and Alarming, Hub Structural Corrosion) 

Support Requested:

From Mission Support Chief Engineer for bringing gray water system back in operation and for propane refill.

From Observatory Team to bring telescope and radiotelescope back in operation

Miscellaneous:


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level: 13.5 Volts (Measured with Voltmeter)

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.75

Propane Tank (Full/0.75/0.5/0.25/Empty): 14%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75


Notes/Comments:

On Mission Support Chief Engineer indications, both GreenHab thermostats have been set to 50 degree last night in order to save propane


Oil:

Kitty Oil Quantity (# of Quarts in storage):

Honey Oil Quantity (# of Quarts in storage):

ATV Oil Quantity (# of Quarts in storage): all 4/4 except Opportunity 0/4 (to be filled)

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 20

Trailer Potable Water Tank Level (inches from bottom): 32

Hab Potable Water Tank Level (inches from bottom): 52

Potable Water Meter Reading: 44043,3


In to GreenHab Meter Reading:


Notes/Comments:

Yesterday night we realized that the pressure pump was keeping periodically (with half a minute frequency) turning on and off. Then we could see water dropping down just over the engineering door.

The heater was on at that moment.

We switched off the heater and after 15 minutes the problems was still persistent

For the night the pump has been stopped, red valve closed and circuit emptied.


Mission Support Chief Engineer informed us that the water heater leaking when on was a known issue.

Indeed, this morning the problem was gone. To be cheched when we will turn the heater on again.


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray water system not functional at the moment.

Waiting for instruction from Mission Support Chief Engineer for bringing it back in operation


Observatory:

Telescope Used? (Yes/No):Yes

Observatory Notes/Comments:

Youval has tried to use the telescope but after some alignment trials he finally got an error message from the telescope control system which has been reported to Peter Detterline and the Observatory Team.


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):yes

Tire Status: good

Odometer at end of day: 123761.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: n

Oil Checked: y

Fuel Consumed: none

Tire Status: to be inflated

Comments:


ATV 2 (Spirit):

Used: n

Oil Checked: y

Fuel Consumed: none

Tire Status: to be inflated

Comments:


ATV 3 (Viking I):

Used: n

Oil Checked: y

Fuel Consumed: none

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: n

Oil Checked: y

Fuel Consumed: none

Tire Status: good

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 66

Actual temperature upstairs (Farenheit): 67

Thermostat setting downstairs (Farenheit):55

Actual temperature downstairs (Farenheit): 61


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:

The setup of the alarming system (F. Carbognani science project) has started. This implied till now only the connection to the network of the PLC controller with address 192.168.1.200.

With this address configuration it was possible to successfully communicate to the PLC via the Modbus over Ethernet protocol.


Science Report (Material Science)

Angeliki Kapoglou Reporting

Corrosion Inspection of the Hab

First inspection of the outside structure was conducted during our first EVA today. Some corroded parts were identified, mostly around the doors and windows, regions that have an occluded geometry in which water and corrosive species can be trapped and also where coating of the aluminum sheets is not present. On the painted areas of the Hab, corrosion was discovered by a change of the color and blisters.Some pores and pinholes were also observed on the outside coating of the aluminum.

This may result in strong local corrosion of the substrate as the coating is not absolutely dense. Photographs of the problematic parts were taken but, since the corrosion initiation cannot always be detected on the surface, more EVAs will follow to define the physical root and types of the present corrosion.


Science Report (Physiology)

Kavya Manyapu Reporting

The crew physiologist K.Manyapu started the physiology experiments. The heart rates, weight, bone mass, muscle mass, lung capacities of the crew have been measured. Three crew members have started on the excercise protocol. The heart rate of the crew performing EVAs before, during and after the EVAs have been measured.

These experiments are ongoing and will be conducted until the end of the mission. The effect of exercise on the health of the crew as well as their performance will be studied.


EVA 1 Report

Kavya Manyapu Reporting

First EVA of the day was commanded by K.Manyapu and was accompanied by A.Kapoglou. The EVAs included the inspection of the science project conducted on the previous days EVA and the inspection of the habitat’s First EVA of the day was commanded by K.Manyapu and was accompanied by A.Kapoglou. The EVAs included the inspection of the science project structural corrosion. The EVA lasted 45 minutes. The cap-com for the EVA was the crew commander F.Carbognani. Because of the duration of the EVA,the crew did face some operational difficulties such as breathing , heat and the reduction of visibility. The glaring Sun was not very helpful in conducting clear inspections. However, the crew managed to get all the data required and were successful in their EVA.


EVA 2 Report

Kavya Manyapu Reporting

EVA 2 followed EVA 1 with a interval of 30 minutes. The EVA 1 crew was commanded by L.Vimericati, the crew Biologist for this EVA. The cap-com of this EVA was F.Carbognani and was assisted to don the spacesuits by the astronauts by Y.Brodsky.  The EVA last for 1.15 hours. The crew traversed a little above half a by foot a mile from the home base in search of rocks containing cyanobacteria and their existence in these extreme environments.  The crew was able to collect samples of the rocks.

The heart rates of the EVA crew was monitored by K.Manyapu to inspect the stress level experienced by the crew while performing EVA tasks.


Lessons Learned:

Due to the high temperatures outside, the crew faced a few operational difficulties.

•       Check the temperature outside the habitat before conducting an EVA.

•       The crew will require to use sunglass protection such that the glaring sun during the morning/afternoon EVAs do not restrict their visibility

•       The microphones need to be more ergonomic to fit the crew ears and be able to communicate properly with the cap-com

•       ATVs must be used to traverse if the EVA is more than 0.2 miles

•       The venting system of the space suits is not very effective in helping the astronauts breath properly


Chef's Report

Y. Brodsky and J. Deaton Reporting

1a) Today was a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration?

Our first time eating tofu, and the first time of the chefs cooking it. It was well received.

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Stir fried tofu with dehydrated onions

Santa Fe chipotle soup

Santa Fe black beans and rice

Dehydrated vegetable medley (boiled)

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

3 servings Santa Fe chipotle soup

2 servings Santa Fe black beans and rice

3 servings Dehydrated vegetable medley (boiled)

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.

Stir fried tofu:

2 boxes tofu

*Vegetable oil

*Dehydrated onions

Heat oil in skillet, dry tofu, dice tofu, add tofu and onions to skillet, stir until golden-brown.

*Quantities added “by feel”.

5)  Lessons learned  (if any)

6) Comments/questions for Kim and Jean, the food study investigators


Journalist Report

Kavya Manyapu Reporting

It has been a busy day for crew 102 with 2 EVAs, implementation of an alarm system for the habitat, start of the exercise protocol for the astronauts and other minor activities. EVA 1 was commanded by for today’s EVAs was the crew commander F.Carbognani.  K.Manyapu checked on the experiment conducted during the previous day EVA for the outreach science program which was on based on dry Ice sublimation insulation by the soil. A.Kapoglou, crew material science engineer took photographs of material corrosion on the MDRS habitat structure which will be carefully studied in the future. The second EVA was commanded by L.Vimericati, accompanied by K.Manyapu and A.Kapoglou. The existence of cyanobacteria in the cracks of the Martian (Analogue) rocks was tested and samples of rock were collected for future testing

K.Manyapu in the morning along with crew mate A.Kapoglou. The Cap-Com in the laboratory conducted by the crew biologist L.Viimericati. Both the EVAs were successful and the astronauts were able to also get some video footage of the EVAs. 

F.Carbognani, assisted by Y. Brodsky, started working on putting together a monitoring and alarm system that measures the temperature, humidity and CO2 levels of the habitat. A provisional set-up will be completed by tonight including the CO2 level alarming. The physiology experiments have also been implemented by the crew physiologist K.Manyapu today. The heart rates, weight and other important physiological parameters have been measured for all the crew today. 3 astronauts have started their exercise protocol as of today which will feed into the effects of exercise on crew health and performance experiments.

Today’s chefs were F.Carbognani, J.Deaton and Y.Brodsky who have made excellent food with the use of the astronaut food that was supplied to the crew. Some of the crew got their first showers today, after 3 days on the Martian land. The crew is looking forward to another exciting day of experiments. Over and Out for today.